3. CP India Ltd has the following capital structure, which it considers optimal:
Debt 25%
Preference Shares 15%
Equity shares 60%
Total 100%
Applicable tax rate for CPIL is 25%. and investors expect earnings and dividends to grow at a constant rate of 9% in the future. Risk free rate of return is 6%, average equity share has expected rate of return of 15%. CPIL’s beta is 1.50. Following terms would
apply to new securities being issued as follows:
1. New preference can be issued at a face value of Rs. 100 per share, dividend and cost of
issuance will be Rs. 8 per share and Rs. 4 per share respectively.
2. Debt will bear an interest rate of 10%.


Calculate
a. Component cost of debt, preference shares and equity shares assuming that CPIL does
not issue any additional equity shares.
b. WACC.
1
Expert's answer
2020-12-23T10:59:31-0500

a. Component cost of debt is:

"0.1\u00d7(1 - 0.25) = 0.075" or 7.5%.

Cost of preference shares is:

"12\/100 = 0.12" or 12%.

and cost of equity shares is:

"Ke = 0.06 + 1.5\u00d7(0.15 - 0.06) = 0.195" or 11.25%.

b. "WACC = 0.195\u00d70.6 + 0.075\u00d70.25 + 0.12\u00d70.15 = 0.1538."